two markets for substitute goods:


Qs1 = 2P1

QS2=-10+2P2


QD1=20-P1+P2

QD2=40-2P2+P1


FIND THE EQUILIBRIUM PRICE AND QUANTITY OF THE TWO GOODS




1
Expert's answer
2022-04-18T17:28:22-0400

"Q_s=Q_d,"

"2P_1=20-P_1+P_2,"

"-10+2P_2=40-2P_2+P_1,"

"P_1=130\/11,"

"P_2=170\/11,"

"Q_1=260\/11,"

"Q_2=230\/11."